# PhysX-3D: Physical-Grounded 3D Asset Generation

### v3

@@ Abstract
3D modeling is moving from virtual to physical. Existing 3D generation primarily emphasizes geometries and textures while neglecting physical-grounded modeling. Consequently, despite the rapid development of 3D generative models, the synthesized 3D assets often overlook rich and important physical properties, hampering their real-world application in physical domains like simulation and embodied AI. As an initial attempt to address this challenge, we propose PhysX-3D, an end-to-end paradigm for physical-grounded 3D asset generation. 1) To bridge the critical gap in physics-annotated 3D datasets, we present PhysXNet - the first physics-grounded 3D dataset systematically annotated across five foundational dimensions: absolute scale, material, affordance, kinematics, and function description. In particular, we devise a scalable human-in-the-loop annotation pipeline based on vision-language models, which enables efficient creation of physics-first assets from raw 3D assets. 2) Furthermore, we propose PhysXGen, a feed-forward framework for physics-grounded image-to-3D asset generation, injecting physical knowledge into the pre-trained 3D structural space. Specifically, PhysXGen employs a dual-branch architecture to explicitly model the latent correlations between 3D structures and physical properties, thereby producing 3D assets with plausible physical predictions while preserving the native geometry quality. Extensive experiments validate the superior performance and promising generalization capability of our framework. All the code, data, and models will be released to facilitate future research in generative physical AI.
@@ 1 Introduction
The creation of diverse and high-quality 3D assets has gained significant prominence in recent years, driven by their expanding applications across gaming, robotics, and embodied simulators. Substantial research efforts have been focused on appearance and geometry only, from high-quality 3D datasets [3, 8, 7, 25], efficient 3D representations, to generative modeling. However, most of them predominantly emphasize structural characteristics while overlooking physical properties inherent to real-world objects. Given the rising demand for physical modeling, understanding, and reasoning in 3D space, we argue that a comprehensive suite for physics-grounded 3D objects is important, from upstream data annotations pipeline to downstream generative modeling.
Beyond purely structural attributes like geometry and appearance, real-world objects intrinsically possess rich physical and semantic characteristics comprising: 1) absolute scale, 2) material, 3) affordance, 4) kinematics, and 5) function descriptions. By integrating these fundamental properties with classical physical principles, we can derive critical dynamic metrics, including gravitational effects, frictional forces, contact region, motion trajectories, and interaction. However, existing datasets/annotation pipelines only offer partial solutions towards physically grounded knowledge in 3D objects that cover the entire spectrum. Recent efforts to support articulated object applications have yielded datasets like PartNet-Mobility [26], which provides 2.7K human-annotated articulated 3D models. Yet, this collection still lacks essential physical descriptors - including dimensional specifications, material composition, and functional affordances - that are crucial for physically accurate simulations and robotics applications.
@@
Most importantly, PhysXNet is built with an efficient, robust, and scalable labeling pipeline. We introduce a human-in-the-loop annotation pipeline to annotate the properties for the existing object-level 3D dataset, i.e., PartNet [17]. The pipeline proceeds in three stages: 1) target visual isolation, in which we render each component via alpha compositing to get the best visual prompts with minimized visual interference. 2) automatic VLM labeling, where a large vision-language model (VLM) to annotate most of the properties; and 3) expert refinement, combining systematic spot-checks with focused human annotation of complex kinematic behaviors. To the best of our knowledge, PhysXNet is the first 3D dataset with abundant properties for each part.
@@ 2 Related Work
Table 1: Comparison of related datasets which can support research in physical 3D generation. While the ABO dataset [6] contains material metadata and keywords, its object-level annotation granularity constrains part-aware applications like robotic manipulation or physical simulation. In contrast, PhysXNet provides part-level annotations.
@@
ShapeNet [3] 51K 2015
PartNet [17] 26K 2019
PartNet-Mobility [26] 2.7K 2020
GAPartNet [9] 1.1K 2022
ABO [6] 7.9K Obj-level Obj-level 2022
OmniObject3D [25] 6K 2023
Objaverse [8] 818K 2023
@@
Due to the time-consuming and expensive in realistic data collection, current large-scale 3D datasets prefer to collect data online [3, 8, 7]. According to the type of 3D data, existing 3D datasets can be divided into synthetic and real-world datasets. To facilitate the development of 3D vision, ShapeNet [3] collects 51,300 CAD models. Building upon it, the PartNet dataset [17] introduces an annotation framework that provides part annotations at significantly finer granularity levels. Furthermore, PartNet-Mobility [26] annotates the kinematic constraints and provides 2.7K articulated 3D objects for 3D vision, especially for embodied AI and robotics. ABO [6] is a high-quality datasets with around 7.9K CAD models with fine-grained geometric and textures. Compared with prior work, it includes the physical dimension, material, and keywords. However, the material information and descriptions focus on object-level, limiting the part-aware applications. Recently, Objaverse [8] has alleviated the scarcity of 3D data. It collects and filters over 800K 3D data. To bridge the gap between synthetic and real data, Omniobject3D [25] provides over 6k high-quality 3D scans. A detailed comparison is shown in Table 1.
@@
As one of the most representative optimization-based method in 3D generation, DreamFusion [18] proposed the SDS loss function. By utilizing the prior knowledge of the 2D diffusion model, it achieves impressive generative performance. Despite various works, optimization-based methods still suffer from the multi-face Janus problem and low optimization efficiency. Recently, benefiting from its impressive efficiency and robustness, feed-foward models [27, 23, 29, 11, 4, 1, 2] have gained more and more attention. However, those methods still focus on geometry and appearance quality, neglecting the physical properties of 3D assets.
@@ 2.3 Articulated and Physical 3D Object Modeling
Articulated object modeling mainly consists of tasks like perception, reconstruction, and generation. Some works try to estimate articulation pose [14] and identify articulation parts [30], while others [22] focus on learn joint parameters from images. In the reconstruction field, existing works try to reconstruct articulated models from RGB [5], RGBD [24], and point cloud [12]. Recently, some methods have tried to generate articulated 3D assets by utilizing a vision-language model [13] or adopting an optimization-based framework [19]. To bridge the critical gap between existing methods with real applications, many works aim to incorporate the physical properties into 3D modeling. Some works try to learn material parameters from videos [32] or images [31], while other methods aim to introduce physical guidance via simulation [16, 28] or physical principles [10].
@@ 3.1 Definition of Physical Properties
As shown in Figure 2, we systematically categorize object properties into three progressive stages: a) Identification - determining the basic nature of the object; b) Function - understanding its potential applications; and c) Operation - detailed usage methodologies. To streamline the annotation process, we posit that the internal composition of a component is homogeneous, exhibiting uniform property invariance throughout its structure. For stages a), we set absolute scaling and material (material name, Young’s modulus, Poisson’s ratio, and density). Besides, for b), we establish functional affordance analysis and function descriptions (basic, functional, and kinematic descriptions). Finally, we use kinematic parameter quantification to represent c). Specifically, we grade the priority of being touched on all available parts to obtain the affordance score for all parts from 1 to 10. We set five possible kinematic types: A. No movement constraints (like water in a bottle), B. Prismatic joints (like a drawer), C. Revolute joints (like a laptop), D. Hinge joint (like a hose in a shower system), or E. Rigid joint and a combined kinematic type: CB. Revolute and Prismatic joints (like a lid of a bottle). Except for A and E, we will annotate the parent, child parts, and detailed kinematic parameters (such as rotation direction, rotation range, and so on). Note that, due to the challenges in precisely quantifying the absolute physical movement range of B, we use the movement range within the 3D coordinate system. Besides, to avoid the unnecessary and meaningless annotation of over-fine-grained parts in PartNet, we merge the tiny parts whose vertices and area are smaller than a pre-defined threshold with their neighboring parts.
@@ 3.2 Human-in-the-loop Annotation Pipeline
Following the establishment of target annotation specifications, we implement a systematic and streamlined semi-automated annotation framework, structured into two distinct operational phases (see Figure 2): 1) Preliminary Data Acquisition and 2) Kinematic Parameter Determination. Specifically, we utilize GPT-4o to obtain the basic information. Besides, to ensure the quality of raw data, a human candidate will check the output of the vision-language model (VLM). For the second phrase, we split it into four subtasks: (2.a) calculate contact region, (2.b) plane fitting, (2.c) candidate generation and selection, and (2.d) kinematic parameters. For all constraint movable parts (kinematic type is not A or E), we will calculate the contact region with the neighboring parts. We first extract point cloud data from the child-parent mesh pair, formally designated as $P_{c}$ and $P_{p}$, respectively. The workflow subsequently calculates Euclidean distance between points in $P_{c}$ and $P_{p}$, followed by spatial filtration that eliminates point pairs failing to meet a predetermined distance threshold. Subsequently, we employ a plane-fitting algorithm. We sample several axes uniformly on the fitted plane as candidates. Note that for kinematic type C, we additionally need to determine the location of the rotation axis. Therefore, we will perform a k-means algorithm in the contact region for type C. After selecting the candidate axis and location, we can finalize the kinematic parameters.
@@ 4.1 Physical 3D VAE Encoding and Decoding
In this subsection, we take the textured mesh output as an example. To reduce the influence caused by the domain gap between geometric and physical latent space, we build a similar physical VAE for property encoding, following [27]. Besides, considering the interdependencies among physical properties, we encode them into a unified latent space. We adopt 4 physical properties: physical scaling (converted by physical dimension) $P_{dim}\in\mathbb{R}^{N\times 1}$, affordance priority $P_{aff}\in\mathbb{R}^{N\times 1}$, density $P_{\rho}\in\mathbb{R}^{N\times 1}$, and kinematic parameters $P_{mov}\in\mathbb{R}^{N\times 11}$ (including child $\mathbb{R}^{N\times 1}$ and parent group index $\mathbb{R}^{N\times 1}$, movement direction $\mathbb{R}^{N\times 3}$, movement location $\mathbb{R}^{N\times 3}$, movement range $\mathbb{R}^{N\times 2}$, and kinematic type $\mathbb{R}^{N\times 1}$), where N is the number of voxel. The physical properties ($P_{phy}\in\mathbb{R}^{N\times 14}$) can be obtained by channel-wise concatenation. For the function descriptions, we adopt the CLIP model [20] to obtain the text embedding. Similarly, the description features ($P_{sem}\in\mathbb{R}^{N\times 768\times 3}$) are formed by concatenating the basic, functional, and kinematic description embeddings. Besides, the structural branch adopts the DINOv2 to extract features. Therefore, the dimensions of structural feature is $P_{aes}\in\mathbb{R}^{N\times 1024}$. For clarification, we denote the pretrain VAE encoder and decoder as $\mathcal{E}_{aes}$ and $\mathcal{D}_{aes}$ while the physical VAE encoder and decoder as $\mathcal{E}_{phy}$ and $\mathcal{D}_{phy}$. The physical latent $P_{plat}\in\mathcal{R}^{N\times 8}$ and structured latent $P_{slat}\in\mathcal{R}^{N\times 8}$ can be formulated as follows:
@@ 4.2 Physical Latent Generation
Following the acquisition of the compressed physical latent representation, we construct a transformer-architecture diffusion model to jointly generate physical and structural attributes. To effectively leverage the inherent correlations between physical properties and structural features while maintaining compatibility with pre-trained components, we implement a dual-branch architecture that integrates structural guidance through residual connections. Specifically, the additional branch from the structural module is fused with the primary physical generation module via learnable skip-connection layers, enabling cross-domain feature interaction. Comprehensive ablation studies quantitatively validate the design rationale through systematic component comparisons. Following [27], we adopt the Conditional Flow Matching (CFM) as the objective of optimization. Therefore, the loss of the geometric branch is formulated:
$\begin{split}\mathcal{L}_{aes}=\mathbb{E}_{t,x_{0},\epsilon}||f(x,t)-(\epsilon-x_{0})||^{2}_{2}\end{split}~{},$ (3)
@@ 5.1 Implementation details
In our experiments, we partition PhysXNet dataset into 24K training samples, 1K validation samples, and 1K test cases. By analyzing the performance on the test cases, we can evaluate the generalizability of our method. During the VAE and diffusion model training, we adopt AdamW with an initial learning rate of $1\times 10^{-4}$ to optimize the models. The inherent correlation between geometric configuration and physical properties in our methodology creates a critical dependency where the structural fidelity of the 3D representation will affect the final generative performance. In this paper, we repurpose the geometry- and appearance-rich structural space of TRELLIS [27] for our task. Our PhysXGen is trained on 8 NVIDIA A100 GPUS. More details about the architecture are released in the supplementary.
@@ 5.2 Evaluation Metrics
Physical properties evaluation. Our framework establishes a multi-property feature space encompassing five core attributes: absolute scale, material, affordance, kinematics, and function descriptors. Note that the kinematics attribute manifests as dual configuration parameters: 1) structural grouping (parent-child part hierarchies) and 2) kinematic parameters. To systematically evaluate these properties, we implement an Euclidean distance-based evaluation paradigm. Specifically, we render the generated object with physical properties from 10 pre-defined views and calculate the MAE of rendered property images and ground-truth.
@@
Methods PSNR $\uparrow$ CD $\downarrow$ F-Score $\uparrow$ Absolute scale $\downarrow$ Material $\downarrow$ Affordance $\downarrow$ Kinematic parameters $\downarrow$ Description $\downarrow$
TRELLIS [27] 24.31 13.2 76.9
TRELLIS + PhysPre 24.31 13.2 76.9 12.46 0.262 0.435 0.589 1.01
PhysXGen 24.53 12.7 77.3 6.63 0.141 0.372 0.479 0.71
@@
Geometry evaluation. For appearance evaluation, we sample 30 random views from a unit sphere to calculate the mean Peak Signal-to-Noise Ratio (PSNR). Besides, to evaluate the quality of geometry, we calculate the standard shape metrics of Chamfer Distance (CD) (×$\times 10^{-3}$) and F-score (FS) (×$\times 10^{-2}$) with thresholds of 0.05.
@@
Dep-VAE Dep-Diff PSNR $\uparrow$ CD $\downarrow$ F-Score $\uparrow$ Absolute scale $\downarrow$ Material $\downarrow$ Affordance $\downarrow$ Kinematic parameters $\downarrow$ Description $\downarrow$
24.31 13.2 76.9 12.46 0.262 0.435 0.589 1.01
24.31 13.2 76.9 11.71 0.202 0.401 0.514 0.81
24.32 12.9 77.0 10.18 0.215 0.391 0.495 0.82
24.53 12.7 77.3 6.63 0.141 0.372 0.479 0.71
Figure 6: Qualitative comparison of different methods. We employ MAE-based similarity assessment between generated results and ground-truth for absolute scale, material, and kinematics, while affordance and function description are evaluated through affordance ranking and cosine similarity score with inputted prompt, respectively.

### v4

@@ Abstract
3D modeling is moving from virtual to physical. Existing 3D generation primarily emphasizes geometries and textures while neglecting physical-grounded modeling. Consequently, despite the rapid development of 3D generative models, the synthesized 3D assets often overlook rich and important physical properties, hampering their real-world application in physical domains like simulation and embodied AI. As an initial attempt to address this challenge, we propose PhysX, an end-to-end paradigm for physical-grounded 3D asset generation. 1) To bridge the critical gap in physics-annotated 3D datasets, we present PhysXNet - the first physics-grounded 3D dataset systematically annotated across five foundational dimensions: absolute scale, material, affordance, kinematics, and function description. In particular, we devise a scalable human-in-the-loop annotation pipeline based on vision-language models, which enables efficient creation of physics-first assets from raw 3D assets. 2) Furthermore, we propose PhysXGen, a feed-forward framework for physics-grounded image-to-3D asset generation, injecting physical knowledge into the pre-trained 3D structural space. Specifically, PhysXGen employs a dual-branch architecture to explicitly model the latent correlations between 3D structures and physical properties, thereby producing 3D assets with plausible physical predictions while preserving the native geometry quality. Extensive experiments validate the superior performance and promising generalization capability of our framework. All the code, data, and models will be released to facilitate future research in generative physical AI.
@@ 1 Introduction
The creation of diverse and high-quality 3D assets has gained significant prominence in recent years, driven by their expanding applications across gaming, robotics, and embodied simulators. Substantial research efforts have been focused on appearance and geometry only, from high-quality 3D datasets shapenet, objaverse, objaversexl, omniobject3d, efficient 3D representations, to generative modeling. However, most of them predominantly emphasize structural characteristics while overlooking physical properties inherent to real-world objects. Given the rising demand for physical modeling, understanding, and reasoning in 3D space, we argue that a comprehensive suite for physics-grounded 3D objects is important, from upstream data annotations pipeline to downstream generative modeling.
Beyond purely structural attributes like geometry and appearance, real-world objects intrinsically possess rich physical and semantic characteristics comprising: 1) absolute scale, 2) material, 3) affordance, 4) kinematics, and 5) function descriptions. By integrating these fundamental properties with classical physical principles, we can derive critical dynamic metrics, including gravitational effects, frictional forces, contact region, motion trajectories, and interaction. However, existing datasets/annotation pipelines only offer partial solutions towards physically grounded knowledge in 3D objects that cover the entire spectrum. Recent efforts to support articulated object applications have yielded datasets like PartNet-Mobility partnetmobility, which provides 2.7K human-annotated articulated 3D models. Yet, this collection still lacks essential physical descriptors - including dimensional specifications, material composition, and functional affordances - that are crucial for physically accurate simulations and robotics applications.
@@
Most importantly, PhysXNet is built with an efficient, robust, and scalable labeling pipeline. We introduce a human-in-the-loop annotation pipeline to annotate the properties for the existing object-level 3D dataset, i.e., PartNet partnet. The pipeline proceeds in three stages: 1) target visual isolation, in which we render each component via alpha compositing to get the best visual prompts with minimized visual interference. 2) automatic VLM labeling, where a large vision‑language model (VLM) to annotate most of the properties; and 3) expert refinement, combining systematic spot‑checks with focused human annotation of complex kinematic behaviors. To the best of our knowledge, PhysXNet is the first 3D dataset with abundant properties for each part.
@@ 2 Related Work
Table 1: Comparison of related datasets which can support research in physical 3D generation. While the ABO dataset abo contains material metadata and keywords, its object-level annotation granularity constrains part-aware applications like robotic manipulation or physical simulation. In contrast, PhysXNet provides part-level annotations.
@@
ShapeNet shapenet 51K 2015
PartNet partnet 26K 2019
PartNet-Mobility partnetmobility 2.7K 2020
GAPartNet gapartnet 1.1K 2022
ABO abo 7.9K Obj-level Obj-level 2022
OmniObject3D omniobject3d 6K 2023
Objaverse objaverse 818K 2023
@@
Due to the time-consuming and expensive in realistic data collection, current large-scale 3D datasets prefer to collect data online shapenet, objaverse, objaversexl. According to the type of 3D data, existing 3D datasets can be divided into synthetic and real-world datasets. To facilitate the development of 3D vision, ShapeNet shapenet collects 51,300 CAD models. Building upon it, the PartNet dataset partnet introduces an annotation framework that provides part annotations at significantly finer granularity levels. Furthermore, PartNet-Mobility partnetmobility annotates the kinematic constraints and provides 2.7K articulated 3D objects for 3D vision, especially for embodied AI and robotics. ABO abo is a high-quality datasets with around 7.9K CAD models with fine-grained geometric and textures. Compared with prior work, it includes the physical dimension, material, and keywords. However, the material information and descriptions focus on object-level, limiting the part-aware applications. Recently, Objaverse objaverse has alleviated the scarcity of 3D data. It collects and filters over 800K 3D data. To bridge the gap between synthetic and real data, Omniobject3D omniobject3d provides over 6k high-quality 3D scans. A detailed comparison is shown in Table 1.
@@
As one of the most representative optimization-based method in 3D generation, DreamFusion poole2022dreamfusion proposed the SDS loss function. By utilizing the prior knowledge of the 2D diffusion model, it achieves impressive generative performance. Despite various works, optimization-based methods still suffer from the multi-face Janus problem and low optimization efficiency. Recently, benefiting from its impressive efficiency and robustness, feed-foward models trellis, lgm, instantmesh, zlabor, 3dtopia, 3dtopiaxl, difftf, difftf++ have gained more and more attention. However, those methods still focus on geometry and appearance quality, neglecting the physical properties of 3D assets.
@@ 2.3 Articulated and Physical 3D Object Modeling
Articulated object modeling mainly consists of tasks like perception, reconstruction, and generation. Some works try to estimate articulation pose liu2022toward and identify articulation parts zengvisual, while others sun2024opdmulti focus on learn joint parameters from images. In the reconstruction field, existing works try to reconstruct articulated models from RGB chen2024urdformer, RGBD weng2024neural, and point cloud hsu2023ditto. Recently, some methods have tried to generate articulated 3D assets by utilizing a vision-language model le2024articulate, cao2025physx or adopting an optimization-based framework qiu2025articulate. To bridge the critical gap between existing methods with real applications, many works aim to incorporate the physical properties into 3D modeling. Some works try to learn material parameters from videos zhong2024reconstruction or images zhai2024physical, while other methods aim to introduce physical guidance via simulation mezghanni2022physical, xie2024physgaussian or physical principles guo2024physically.
@@ 3.1 Definition of Physical Properties
As shown in Figure 2, we systematically categorize object properties into three progressive stages: a) Identification - determining the basic nature of the object; b) Function - understanding its potential applications; and c) Operation - detailed usage methodologies. To streamline the annotation process, we posit that the internal composition of a component is homogeneous, exhibiting uniform property invariance throughout its structure. For stages a), we set absolute scaling and material (material name, Young’s modulus, Poisson’s ratio, and density). Besides, for b), we establish functional affordance analysis and function descriptions (basic, functional, and kinematic descriptions). Finally, we use kinematic parameter quantification to represent c). Specifically, we grade the priority of being touched on all available parts to obtain the affordance score for all parts from 1 to 10. We set five possible kinematic types: A. No movement constraints (like water in a bottle), B. Prismatic joints (like a drawer), C. Revolute joints (like a laptop), D. Hinge joint (like a hose in a shower system), or E. Rigid joint and a combined kinematic type: CB. Revolute and Prismatic joints (like a lid of a bottle). Except for A and E, we will annotate the parent, child parts, and detailed kinematic parameters (such as rotation direction, rotation range, and so on). Note that, due to the challenges in precisely quantifying the absolute physical movement range of B, we use the movement range within the 3D coordinate system. Besides, to avoid the unnecessary and meaningless annotation of over-fine-grained parts in PartNet, we merge the tiny parts whose vertices and area are smaller than a pre-defined threshold with their neighboring parts. We manually refine the results of the merging process to ensure that the merged outputs are reasonable and consistent.
@@ 3.2 Human-in-the-loop Annotation Pipeline
Following the establishment of target annotation specifications, we implement a systematic and streamlined semi-automated annotation framework, structured into two distinct operational phases (see Figure 2): 1) Preliminary Data Acquisition and 2) Kinematic Parameter Determination. Specifically, we utilize GPT-4o to obtain the basic information. Besides, to ensure the quality of raw data, a human candidate will check and refine the output of the vision-language model (VLM).
For the second phrase, we split it into four subtasks: (2.a) calculate contact region, (2.b) plane fitting, (2.c) candidate generation and selection, and (2.d) kinematic parameters. Note that (2.c) and (2.d) are accomplished by human candidate. For all constraint movable parts (kinematic type is not A or E), we will calculate the contact region with the neighboring parts. We first extract point cloud data from the child-parent mesh pair, formally designated as $P_{c}$ and $P_{p}$, respectively. The workflow subsequently calculates Euclidean distance between points in $P_{c}$ and $P_{p}$, followed by spatial filtration that eliminates point pairs failing to meet a predetermined distance threshold. Subsequently, we employ a plane-fitting algorithm. We sample several axes uniformly on the fitted plane as candidates. Note that for kinematic type C, we additionally need to determine the location of the rotation axis. Therefore, we will perform a k-means algorithm in the contact region for type C to generate several candidates. After selecting the candidate location, we can finalize the kinematic parameters.
@@ 4.1 Physical 3D VAE Encoding and Decoding
In this subsection, we take the textured mesh output as an example. To reduce the influence caused by the domain gap between geometric and physical latent space, we build a similar physical VAE for property encoding, following trellis. Besides, considering the interdependencies among physical properties, we encode them into a unified latent space. We adopt 4 physical properties: physical scaling (converted by physical dimension) $P_{dim}\in\mathbb{R}^{N\times 1}$, affordance priority $P_{aff}\in\mathbb{R}^{N\times 1}$, density $P_{\rho}\in\mathbb{R}^{N\times 1}$, and kinematic parameters $P_{mov}\in\mathbb{R}^{N\times 11}$ (including child $\mathbb{R}^{N\times 1}$ and parent group index $\mathbb{R}^{N\times 1}$, movement direction $\mathbb{R}^{N\times 3}$, movement location $\mathbb{R}^{N\times 3}$, movement range $\mathbb{R}^{N\times 2}$, and kinematic type $\mathbb{R}^{N\times 1}$), where N is the number of voxel. The physical properties ($P_{phy}\in\mathbb{R}^{N\times 14}$) can be obtained by channel-wise concatenation. For the function descriptions, we adopt the CLIP model radford2021learning to obtain the text embedding. Similarly, the description features ($P_{sem}\in\mathbb{R}^{N\times 768\times 3}$) are formed by concatenating the basic, functional, and kinematic description embeddings. Besides, the structural branch adopts the DINOv2 to extract features. Therefore, the dimensions of structural feature is $P_{aes}\in\mathbb{R}^{N\times 1024}$. For clarification, we denote the pretrain VAE encoder and decoder as $\mathcal{E}_{aes}$ and $\mathcal{D}_{aes}$ while the physical VAE encoder and decoder as $\mathcal{E}_{phy}$ and $\mathcal{D}_{phy}$. The physical latent $P_{plat}\in\mathcal{R}^{N\times 8}$ and structured latent $P_{slat}\in\mathcal{R}^{N\times 8}$ can be formulated as follows:
@@ 4.2 Physical Latent Generation
Following the acquisition of the compressed physical latent representation, we construct a transformer-architecture diffusion model to jointly generate physical and structural attributes. To effectively leverage the inherent correlations between physical properties and structural features while maintaining compatibility with pre-trained components, we implement a dual-branch architecture that integrates structural guidance through residual connections. Specifically, the additional branch from the structural module is fused with the primary physical generation module via learnable skip-connection layers, enabling cross-domain feature interaction. Comprehensive ablation studies quantitatively validate the design rationale through systematic component comparisons. Following trellis, we adopt the Conditional Flow Matching (CFM) as the objective of optimization. Therefore, the loss of the geometric branch is formulated:
[…]
@@ 5.1 Implementation details
In our experiments, we partition PhysXNet dataset into 24K training samples, 1K validation samples, and 1K test cases. By analyzing the performance on the test cases, we can evaluate the generalizability of our method. During the VAE and diffusion model training, we adopt AdamW with an initial learning rate of $1\times 10^{-4}$ to optimize the models. The inherent correlation between geometric configuration and physical properties in our methodology creates a critical dependency where the structural fidelity of the 3D representation will affect the final generative performance. In this paper, we repurpose the geometry- and appearance-rich structural space of TRELLIS trellis for our task. Our PhysXGen is trained on 8 NVIDIA A100 GPUS. More details about the architecture are released in the supplementary.
@@ 5.2 Evaluation Metrics
Physical properties evaluation. Our framework establishes a multi-property feature space encompassing five core attributes: absolute scale, material, affordance, kinematics, and function descriptors. Note that the kinematics attribute manifests as dual configuration parameters: 1) structural grouping (parent-child part hierarchies) and 2) kinematic parameters. Specifically, we evaluate absolute scale using Euclidean distance, density and affordance images via Peak Signal-to-Noise Ratio (PSNR), kinematics with instantiation distance lei2023nap, and functional description through PSNR on cosine similarity score maps.
@@
Methods Geometry Absolute scale $\downarrow$ Material $\uparrow$ Affordance $\uparrow$ Kinematic parameters Description $\uparrow$
PSNR $\uparrow$ CD $\downarrow$ F-Score $\uparrow$     COV $\uparrow$   MMD $\downarrow$
TRELLIS trellis 24.31 13.2 76.9
TRELLIS + PhysPre 24.31 13.2 76.9 13.21 8.63 7.23 0.24 0.12 6.55
PhysXGen 24.53 12.7 77.3 7.24 13.01 11.30 0.33 0.08 10.11
@@
Geometry evaluation. For appearance evaluation, we sample 30 random views from a unit sphere to calculate the mean PSNR. Besides, to evaluate the quality of geometry, we calculate the standard shape metrics of Chamfer Distance (CD) ($\times 10^{-3}$) and F-score (FS) ($\times 10^{-2}$) with thresholds of 0.05.
@@
Dep-VAE Dep-Diff Geometry Absolute scale $\downarrow$ Material $\uparrow$ Affordance $\uparrow$ Kinematic parameters Description $\uparrow$
PSNR $\uparrow$ CD $\downarrow$ F-Score $\uparrow$     COV $\uparrow$   MMD $\downarrow$
24.31 13.2 76.9 13.21 8.63 7.23 0.24 0.12 6.55
24.31 13.2 76.9 12.01 10.69 8.95 0.26 0.11 7.71
24.32 12.9 77.0 10.57 9.86 9.32 0.28 0.11 7.54
24.53 12.7 77.3 7.24 13.01 11.30 0.33 0.08 10.11
Figure 6: Qualitative comparison of different methods. Compared with our baseline, PhysXGen achieves significant improvements, clearly demonstrating its strong performance in physics-grounded 3D generation.